OpenStudy (anonymous):

y varies inversely as twice x. When x = 2, y = 3. Find y when x = 6. forrreal doe help lol

OpenStudy (anonymous):

@jim_thompson5910 what you wrote is a direct relationship instead of inverse.

jimthompson5910 (jim_thompson5910):

oh wait, i misread...oops

OpenStudy (anonymous):

Good catch, juan

jimthompson5910 (jim_thompson5910):

y varies inversely as twice x means y = k/(2x)

jimthompson5910 (jim_thompson5910):

Plug in x = 2 and y = 3 and solve for k y = k/(2x) 3 = k/(2*2) ...keep going...